mass m moves-eo ar ho wn n the veric ween the ba on the rai negiai b bo ansim ittoCompres. At. pia F the fod mood mokeyt ar and starts the emat hov opp osite 」ap ue maximum compres on p7,the sprin! (tkweot He egmeut DE) 긺 e magh
Answer #1

considering conservation of energy

total energy at releasing point at A = energy at C + fricitional energy loss in section CD + spring energy stored.

mgH = mgh + u mg h + 0.5* k x^2

2 mgh - mgh - 0.25 mgh = 0.5 k x^2

0.75 mgh = 0.5 k x^2

(3/4) mgh = (1/2) k x^2

x^2 = (3/2) mgh / k

x = sqrt ( 3 mgh / 2k)

option (f) is correct

========

do comment in case any doubt, will reply for sure.. Goodluck